# Zilber’s Trichotomy in Hausdorff Geometric Structures

### v1

@@ 5.2. Generic Non-trivial $r$-Hypersurfaces
The next three subsections complete the proof of Theorem 4.10, using Proposition 5.4 as the main tool. They are analogous to subsections 8.4-8.6 of [Cas23]. These remaining cases are essentially word-for-word copies of their analogs in [Cas23], so we will be brief. As in [Cas23], we use the following:
@@
As in [Cas23], we have the following properties:
@@ 5.3. Stationary Non-trivial $r$-Hypersurfaces
We next prove Theorem 4.10 for all stationary non-trivial hypersurfaces and all coordinate-wise generic points, without the added genericity assumptions of Proposition 5.29. As in [Cas23, §8.5], the trick is to use a sequence of independent plane curves to ‘translate’ the initial setup into a more generic one. That this is possible is provided by the local homeomorphism property of Definition 3.1.
@@
As described above, the first step is to ‘translate’ $X$ and $\hat{x}$ along a sequence of ‘independent plane curves’. The following is straightforward and independent of the topology, and is proven in [Cas23] using pure rank and dimension computations:
@@
See Lemmas 8.56 and 8.59 of [Cas23]. In particular, (5) is by the proof of Lemma 8.59. ∎
@@
The rest of the proof of Proposition 5.30 is identical to the analogous proof in[Cas23], so we just give a brief outline.
@@ 6.1. Weakly Generic Intersections
    Suppose $t\in T$ and $u\in U$ are generic over $A$, and $x$ is generic in both $C_{t}$ over $At$, and $D_{u}$ over $Au$. Then $(x,t,u)$ is a weakly generic $(\mathcal{C},\mathcal{D})$-intersection.
@@
    There is a tuple $(x,t,u)$ satisfying the assumptions of (1) above. In particular, there is a weakly generic $(\mathcal{C},\mathcal{D})$-intersection.
@@
    Every weakly generic $(\mathcal{C},\mathcal{D})$-intersection is strongly approximable. In particular, there is a strongly generic $(\mathcal{C},\mathcal{D})$-intersection.
@@
    Everything is clear except the genericity of $x\in M^{2}$. But by Lemma 4.13, $x$ is coherent (because $t$ is coherent and $x$ is coherent over $t$). So it suffices to show that $x$ is $\mathcal{M}$-generic in $M^{2}$. This is well known easy (see [Cas23, Lemma 2.40]).
@@
    Let $I$ be the set of $(x,t,u)$ with $x\in C_{t}\cap D_{u}$. Fix $(x,t,u)\in I$ generic. Then $(x,t,u)$ is such a tuple (this also follows from [Cas23, Lemma 2.40]).
@@
    Let ${{}_{x}}D$ be the set of $u^{\prime}$ with $x\in D_{u^{\prime}}$. Then $u$ is generic in ${{}_{x}}D$ over $x$. By the Baire category axiom, there is some $u^{\prime}\in V$ which is generic in ${{}_{x}}D$ over $tx$. It follows easily that $(x,t,u^{\prime})$ is a strongly generic $(\mathcal{C},\mathcal{D})$-intersection (again, using [Cas23, Lemma 2.40]).
@@ 6.3. Detection of Generic Multiple Intersections
Condition (3) does not matter to us. We include it for consistency with [Cas23].
@@
We now prove, in analogy to [Cas23, Theorem 9.1]:
@@ 6.4. Purity of Ramification and the Dimension of $M$
Our next goal is to show that in contexts where purity of ramification applies, Theorem 6.10 is enough to show that $\dim(M)=1$. This is analogous to section 6 of [Cas23]. In the abstract setting, the notion of purity of ramification is more difficult to state – the issue being that we need to identify an appropriate class of maps which could reasonably be expected to have the relevant property (see Remark 6.17). To remedy this, we will use the terminology of weakly generic intersections developed in the previous subsection, with one addition:
@@ 7. Definable Slopes and Interpreting a Group
Once we have our definable ‘Taylor expansions’, we then build a group in a similar way to previous trichotomy papers (e.g. [HS24] and [Cas23]) – by considering a collection of ‘slopes’ equipped with a composition operation. As in [Cas23], we find it most convenient to work with groupoids rather than groups, because this allows us to work solely with generic points of curves. In the end, this is still sufficient to recover a group, by Hrushovski’s group configuration theorem (see Fact 7.3).
@@ 7.2. Definable Slopes
We say that $(\mathcal{K},\tau)$ has definable slopes if there are groupoids $\{\mathcal{I}A_{n}\}_{n\in\mathbb{Z}^{+}}$, and covariant functors $T_{m,n}:\mathcal{IA}_{m}\rightarrow\mathcal{IA}_{n}$ for all $n<m\in\mathbb{Z}^{+}\cup\{\infty\}$, such that:
@@
    (Coordinatization of Slopes) Suppose $f:x\rightarrow y$ is a morphism in $\mathcal{IA}_{n+1}$ for some $n<\infty$, and let $g=T_{n+1,n}(f)$. Then $g$ is $\mathcal{K}(f)$-definable, and $\dim(f/g)\leq 1$.
@@ 7.3. Coherent Slopes and Codes
Let $n$ be a positive integer. By a coherent $n$-slope configuration, we mean a 5-tuple $(x,y,C,c,\alpha)$ such that:
@@ 7.6. Detecting Tangency and Finding Codes
We now turn to the task of coding $n_{0}$-slopes. The following is an analog of the notion of ‘detecting generic non-transversalities’ from [Cas23]:
@@
We will show that if $\mathcal{M}$ detects tangency, every coherent $n_{0}$-slope has a code, and thus by Proposition 7.33, $\mathcal{M}$-interprets a strongly minimal group. In [Cas23], one has a clear construction of codes: indeed, in ACF0, a rank 2 family of plane curves realizes every non-algebraic coherent $n_{0}$-slope finitely many times, and one can code a slope using its finite set of occurrences in any such family. However, in the current setting, there is no reason that a non-algebraic coherent slope must occur in a rank 2 family. So we might have an infinite set of occurrences in a larger family, any two of which are $\mathcal{M}$-dependent by assumption, and we want to ‘code’ this infinite set. The proper tool for doing such a thing is canonical bases. As it turns out, we will code each slope using the canonical base of the dependence between two independent occurrences of that slope in a family.
@@ 7.7. Tangent and Multiple Intersections
    If $n\in\mathbb{Z}^{+}$ and $V\subset K^{2}\times T\times U$ is any open set, we say that $\mathcal{X}$ and $\mathcal{Y}$ have $n$-branching in $V$ over $A$ if for all strongly generic $(\mathcal{X},\mathcal{Y})$-intersections $(x,y,t,u)$ over $A$, if $(x,y,t,u)\in V$ then $\alpha_{xy}^{n}(X_{t})\neq\alpha_{xy}^{n}(Y_{u})$.
@@
    A generic $(\mathcal{X},\mathcal{Y})$-tangency over $A$ is a tuple $(x,y,t,u,n)$ such that $(x,y,t,u)$ is a weakly generic $(\mathcal{X},\mathcal{Y})$-intersection, $n\in\mathbb{Z}^{+}$, $\alpha_{xy}^{n}(X_{t})=\alpha_{xy}^{n}(Y_{u})$, and $X$ and $Y$ have $n$-branching in a neighborhood of $(x,y,t,u)$.
@@
Assume that $(\mathcal{K},\tau,\{\mathcal{IA}n\},\{T_{mn}\})$ satisfies TIMI, and $\mathcal{M}$ detects multiple intersections. Then $\mathcal{M}$ detects tangency. In particular, $\mathcal{M}$ interprets a strongly minimal group.
@@
Let us proceed. Let $d_{i}=\operatorname{rk}(c_{i})$ for $i=1,2$. By Lemma 7.21, $(x_{0},y_{0})$ is generic in $M^{2}$, and each $\dim(c_{i}/x_{0}y_{0})=d_{i}-1$. Thus $d_{i}\geq 1$. Now by Lemma 6.8, we can realize each $C_{i}$ on a generic curve in a standard family (see Definition 2.4 and Lemma 6.8). That is, we can find standard families of non-trivial plane curves $\mathcal{Z}=\{Z_{t}:t\in T\}$ and $\mathcal{W}=\{W_{u}:u\in U\}$ over $\emptyset$, and generic $t_{0}\in T$ and $u_{0}\in U$, so that $C_{1}$ is almost contained in $Z_{t_{0}}$ and $C_{2}$ is almost contained in $W_{u_{0}}$. For each $t\in T$, let $X_{t}=\rho(Z_{t})$. Similarly, for each $u\in U$, let $Y_{u}=\rho(W_{u})$. Then $\mathcal{X}=\{X_{t}:t\in T\}$ and $\mathcal{Y}=\{Y_{u}:u\in U\}$ are $\mathcal{K}(\emptyset)$-definable families of correspondences in $K$. Note that $\dim(T)=d_{1}$ and $\dim(U)=d_{2}$.
We want to apply TIMI to $\mathcal{X}$ and $\mathcal{Y}$, at the tuple $(\rho(x_{0}),\rho(y_{0}),t_{0},u_{0},n)$. First, note that $(x_{0},y_{0},t_{0},u_{0})$ is a weakly generic $(\mathcal{Z},\mathcal{W})$-intersection over $\emptyset$. It follows easily that $(\rho(x_{0}),\rho(y_{0}),t_{0},u_{0})$ is a weakly generic $(\mathcal{X},\mathcal{Y})$-intersection over $\emptyset$.
@@ 9.1. Visceral structrues
In [DG22] Dolich and Goodrick introduce visceral structures as a common generalisation of o-minimality and P-minimality, that – when restricted to the dp-minimal setting – is very similar to the tame uniform structures of Simon and Walsberg ([SW19]) called SW-uniformities in [HHP22]. This formalism also covers the context of 1-h-minimal structures (as pointed out in [CHRV21, Example 2.2.2]). It follows immediately from the results of [DG22] that certain strengthenings of visceral structures considered in that paper are Hausdorff geometric structures.
@@
Hensel minimal structures, introduced in [CHRK22] (for residue characteristic 00 and in [CHRV21] for positive residue characteristic) cover all pure non-trivially valued henselian fields (of characteristic 00) as well as many interesting expansions (covering the $V$-minimal algebraically closed valued fields of [HK06] and the $T$-convex expansions of power-bounded o-minimal fields, [vdDL95]). One sub-class of Hensel minimal structures of special interest (covering all the above examples) is the class of 1-h-minimal fields. As already mentioned, 1-h-minimal structures are visceral, and it is not hard to verify that they are geometric structures (e.g., [AH24, Proposition 2.11]). While 1-h-minimal fields do not necessarily have DFC, every 1-h-minimal field has a 1-h-minimal expansion which does ([CHRV21, Proposition 3.2.3]). Thus, we have:
@@
In both $\mathcal{K}$ and $\mathcal{K}^{\prime}$, the topology is the valuation topology, and it is definable. Since the dimension in both structures is determined by the topology (see, e.g., [AH24, Proposition 2.11]), it follows that for a $\mathcal{K}$-definable set $X$ we have $\dim_{\mathcal{K}}(X)=\dim_{\mathcal{K}^{\prime}}(X)$. Now consider $X$ a $\mathcal{K}$-definable set. Then $\mathrm{Fr}(X)$ is also $\mathcal{K}$-definable, and by the frontier inequality in $\mathcal{K}^{\prime}$ we have that
@@
First, we have to define a notion of smoothness meeting Definition 3.27. We declare the smooth locus of a definable set $X$ to be the set of points $x\in X$ where $X$ is, locally near $x$, a $\dim(X)$-dimensional weak strictly differentiable manifold (see [AH24, Definition 5.3]). By visceral cell decomposition, every definable set is generically locally a topological manifold, and by Proposition 3.12 loc. cit. it is, in fact, generically, a strictly differentiable manifold. This gives all the properties of Definition 3.27 except maybe (5), the preservation of smoothness under preimages of “nice enough” projections. But recall (see the remark following Definition 3.27) that this clause is a generalization of, and would follow from, the ‘submersion theorem’ of differential geometry. Moreover, there is a submersion theorem for 1-h-minimal field ([AH24, Proposition 4.10]). So we simply apply that proposition exactly as in the earlier remark.
@@
We use the usual notion of a tangent space associated with the notion of (weak) strictly differentiable manifolds ([AH24, Definition 5.4]). All properties (1)-(7) of the definition are automatic, and we don’t dwell on them. Clause (8) of the definition, asserting that a definable morphism inducing an isomorphism on the level of tangent spaces is open (locally, near the point in question) is an immediate consequence of the Inverse Mapping Theorem ([AH24, Proposition 4.4]). So it only remains to check that if $f:(X,x)\to(Y,y)$ is such that $x\in X$ and $y\in Y$ are generic, then the induced map between the tangent spaces is surjective. This, in turn, is an immediate result of the corresponding version of Sard’s Theorem [AH24, Proposition 5.21], asserting that the set of singular values of $f$ is non-generic.

### v2

@@ 5.2. Generic Non-trivial $r$-Hypersurfaces
The next three subsections complete the proof of Theorem 4.10, using Proposition 5.4 as the main tool. They are analogous to subsections 8.4-8.6 of [Cas24b]. These remaining cases are essentially word-for-word copies of their analogs in [Cas24b], so we will be brief. As in [Cas24b], we use the following:
@@
As in [Cas24b], we have the following properties:
@@ 5.3. Stationary Non-trivial $r$-Hypersurfaces
We next prove Theorem 4.10 for all stationary non-trivial hypersurfaces and all coordinate-wise generic points, without the added genericity assumptions of Proposition 5.29. As in [Cas24b, §8.5], the trick is to use a sequence of independent plane curves to ‘translate’ the initial setup into a more generic one. That this is possible is provided by the local homeomorphism property of Definition 3.1.
@@
As described above, the first step is to ‘translate’ $X$ and $\hat{x}$ along a sequence of ‘independent plane curves’. The following is straightforward and independent of the topology, and is proven in [Cas24b] using pure rank and dimension computations:
@@
See Lemmas 8.56 and 8.59 of [Cas24b]. In particular, (5) is by the proof of Lemma 8.59. ∎
@@
The rest of the proof of Proposition 5.30 is identical to the analogous proof in[Cas24b], so we just give a brief outline.
@@ 6.1. Weakly Generic Intersections
    Suppose $t\in T$ and $u\in U$ are generic over $A$, and $x$ is generic in both $C_{t}$ over $At$, and $D_{u}$ over $Au$. Then $(x,t,u)$ is a weakly generic $(\mathcal{C},\mathcal{D})$-intersection over $A$.
@@
    There is a tuple $(x,t,u)$ satisfying the assumptions of (1) above. In particular, there is a weakly generic $(\mathcal{C},\mathcal{D})$-intersection over $A$.
@@
    Every weakly generic $(\mathcal{C},\mathcal{D})$-intersection over $A$ is strongly approximable over $A$. In particular, there is a strongly generic $(\mathcal{C},\mathcal{D})$-intersection over $A$.
@@
    Everything is clear except the genericity of $x\in M^{2}$. But by Lemma 4.13, $x$ is coherent (because $t$ is coherent and $x$ is coherent over $t$). So it suffices to show that $x$ is $\mathcal{M}$-generic in $M^{2}$. This is well known easy (see [Cas24b, Lemma 2.40]).
@@
    Let $I$ be the set of $(x,t,u)$ with $x\in C_{t}\cap D_{u}$. Fix $(x,t,u)\in I$ generic. Then $(x,t,u)$ is such a tuple (this also follows from [Cas24b, Lemma 2.40]).
@@
    Let ${{}_{x}}D$ be the set of $u^{\prime}$ with $x\in D_{u^{\prime}}$. Then $u$ is generic in ${{}_{x}}D$ over $x$. By the Baire category axiom, there is some $u^{\prime}\in V$ which is generic in ${{}_{x}}D$ over $tx$. It follows easily that $(x,t,u^{\prime})$ is a strongly generic $(\mathcal{C},\mathcal{D})$-intersection (again, using [Cas24b, Lemma 2.40]).
@@ 6.3. Detection of Generic Multiple Intersections
Condition (3) does not matter to us. We include it for consistency with [Cas24b].
@@
We now prove, in analogy to [Cas24b, Theorem 9.1]:
@@ 6.4. Purity of Ramification and the Dimension of $M$
Our next goal is to show that in contexts where purity of ramification applies, Theorem 6.10 is enough to show that $\dim(M)=1$. This is analogous to section 6 of [Cas24b]. In the abstract setting, the notion of purity of ramification is more difficult to state – the issue being that we need to identify an appropriate class of maps which could reasonably be expected to have the relevant property (see Remark 6.17). To remedy this, we will use the terminology of weakly generic intersections developed in the previous subsection, with one addition:
@@ 7. Definable Slopes and Interpreting a Group
Once we have our definable ‘Taylor expansions’, we then build a group in a similar way to previous trichotomy papers (e.g. [HS25] and [Cas24b]) – by considering a collection of ‘slopes’ equipped with a composition operation. As in [Cas24b], we find it most convenient to work with groupoids rather than groups, because this allows us to work solely with generic points of curves. In the end, this is still sufficient to recover a group, by Hrushovski’s group configuration theorem (see Fact 7.3).
@@ 7.2. Definable Slopes
We say that $(\mathcal{K},\tau)$ has definable slopes if there are groupoids $\{\mathcal{I}A_{n}\}_{n\in\mathbb{N}}$, and covariant functors $T_{m,n}:\mathcal{IA}_{m}\rightarrow\mathcal{IA}_{n}$ for all $n<m\in\mathbb{N}$, such that:
@@
    (Coordinatization of Slopes) Suppose $f:x\rightarrow y$ is a morphism in $\mathcal{IA}_{n}$ for some $n<\infty$. If $n=0$, then $f$ is $\mathcal{K}$-interdefinable with $xy$. If $n>1$, and $g=T_{n,n-1}(f)$, then $g$ is $\mathcal{K}(f)$-definable and $\dim(f/g)\leq 1$.
@@ 7.3. Coherent Slopes and Codes
Let $n$ be a non-negative integer. By a coherent $n$-slope configuration, we mean a 5-tuple $(x,y,C,c,\alpha)$ such that:
@@
By Definition 7.10(6), note that every coherent 0-slope is algebraic. So necessarily $n_{0}\geq 1$, and thus every coherent $n_{0}$-slope has an $n_{0}-1$-th truncation. We use this freely below.
@@ 7.6. Detecting Tangency and Finding Codes
We now turn to the task of coding $n_{0}$-slopes. The following is an analog of the notion of ‘detecting generic non-transversalities’ from [Cas24b]:
@@
We will show that if $\mathcal{M}$ detects tangency, every coherent $n_{0}$-slope has a code, and thus by Proposition 7.33, $\mathcal{M}$-interprets a strongly minimal group. In [Cas24b], one has a clear construction of codes: indeed, in ACF0, a rank 2 family of plane curves realizes every non-algebraic coherent $n_{0}$-slope finitely many times, and one can code a slope using its finite set of occurrences in any such family. However, in the current setting, there is no reason that a non-algebraic coherent slope must occur in a rank 2 family. So we might have an infinite set of occurrences in a larger family, any two of which are $\mathcal{M}$-dependent by assumption, and we want to ‘code’ this infinite set. The proper tool for doing such a thing is canonical bases. As it turns out, we will code each slope using the canonical base of the dependence between two independent occurrences of that slope in a family.
@@ 7.7. Tangent and Multiple Intersections
    If $n\in\mathbb{N}$ and $V\subset K^{2}\times T\times U$ is any open set, we say that $\mathcal{X}$ and $\mathcal{Y}$ have $n$-branching in $V$ over $A$ if for all strongly generic $(\mathcal{X},\mathcal{Y})$-intersections $(x,y,t,u)$ over $A$, if $(x,y,t,u)\in V$ then $\alpha_{xy}^{n}(X_{t})\neq\alpha_{xy}^{n}(Y_{u})$.
@@
    A generic $(\mathcal{X},\mathcal{Y})$-tangency over $A$ is a tuple $(x,y,t,u,n)$ such that $(x,y,t,u)$ is a strongly approximable weakly generic $(\mathcal{X},\mathcal{Y})$-intersection, $n\in\mathbb{N}$, $\alpha_{xy}^{n}(X_{t})=\alpha_{xy}^{n}(Y_{u})$, and $X$ and $Y$ have $n$-branching in a neighborhood of $(x,y,t,u)$.
@@
Assume that $(\mathcal{K},\tau,\{\mathcal{IA}_{n}\},\{T_{mn}\})$ satisfies TIMI, and $\mathcal{M}$ detects multiple intersections. Then $\mathcal{M}$ detects tangency. In particular, $\mathcal{M}$ interprets a strongly minimal group.
@@
Let us proceed. Let $d_{i}=\operatorname{rk}(c_{i})$ for $i=1,2$. By Lemma 7.21, $(x_{0},y_{0})$ is generic in $M^{2}$, and each $\dim(c_{i}/x_{0}y_{0})=d_{i}-1$. It follows that $d_{i}\geq 2$: indeed, otherwise $\dim(c_{i}/x_{0}y_{0})=0$, implying that $\dim(\alpha/x_{0}y_{0})=0$ and contradicting that $\alpha$ is non-algebraic.
Now by Lemma 6.8, we can realize each $C_{i}$ on a generic curve in a standard family (see Definition 2.4 and Lemma 6.8). That is, we can find standard families of non-trivial plane curves $\mathcal{Z}=\{Z_{t}:t\in T\}$ and $\mathcal{W}=\{W_{u}:u\in U\}$ over $\emptyset$, and generic $t_{0}\in T$ and $u_{0}\in U$, so that $C_{1}$ is almost contained in $Z_{t_{0}}$ and $C_{2}$ is almost contained in $W_{u_{0}}$. For each $t\in T$, let $X_{t}=\rho(Z_{t})$. Similarly, for each $u\in U$, let $Y_{u}=\rho(W_{u})$. Then $\mathcal{X}=\{X_{t}:t\in T\}$ and $\mathcal{Y}=\{Y_{u}:u\in U\}$ are $\mathcal{K}(\emptyset)$-definable families of correspondences in $K$. Note that $\dim(T)=d_{1}$ and $\dim(U)=d_{2}$.
We want to apply TIMI to $\mathcal{X}$ and $\mathcal{Y}$, at the tuple $(\rho(x_{0}),\rho(y_{0}),t_{0},u_{0},n)$. First, note that $(x_{0},y_{0},t_{0},u_{0})$ is a weakly generic $(\mathcal{Z},\mathcal{W})$-intersection over $\emptyset$ – and by Lemma 6.3(3) (recalling that $d_{i}\geq 2$), $(x_{0},y_{0},t_{0},u_{0})$ is also strongly approximable over $\emptyset$. It follows easily that $(\rho(x_{0}),\rho(y_{0}),t_{0},u_{0})$ is a strongly approximable weakly generic $(\mathcal{X},\mathcal{Y})$-intersection over $\emptyset$.
@@ 9.1. Visceral structrues
In [DG22] Dolich and Goodrick introduce visceral structures as a common generalisation of o-minimality and P-minimality, that – when restricted to the dp-minimal setting – is very similar to the tame uniform structures of Simon and Walsberg ([SW19]) called SW-uniformities in [HHP22]. This formalism also covers the context of 1-h-minimal structures (as pointed out in [CHRKV23, Example 2.2.2]). It follows immediately from the results of [DG22] that certain strengthenings of visceral structures considered in that paper are Hausdorff geometric structures.
@@
Hensel minimal structures, introduced in [CHRK22] (for residue characteristic 00 and in [CHRKV23] for positive residue characteristic) cover all pure non-trivially valued henselian fields (of characteristic 00) as well as many interesting expansions (covering the $V$-minimal algebraically closed valued fields of [HK06] and the $T$-convex expansions of power-bounded o-minimal fields, [vdDL95]). One sub-class of Hensel minimal structures of special interest (covering all the above examples) is the class of 1-h-minimal fields. As already mentioned, 1-h-minimal structures are visceral, and it is not hard to verify that they are geometric structures (e.g., [ALH25, Proposition 2.11]). While 1-h-minimal fields do not necessarily have DFC, every 1-h-minimal field has a 1-h-minimal expansion which does ([CHRKV23, Proposition 3.2.3]). Thus, we have:
@@
In both $\mathcal{K}$ and $\mathcal{K}^{\prime}$, the topology is the valuation topology, and it is definable. Since the dimension in both structures is determined by the topology (see, e.g., [ALH25, Proposition 2.11]), it follows that for a $\mathcal{K}$-definable set $X$ we have $\dim_{\mathcal{K}}(X)=\dim_{\mathcal{K}^{\prime}}(X)$. Now consider $X$ a $\mathcal{K}$-definable set. Then $\mathrm{Fr}(X)$ is also $\mathcal{K}$-definable, and by the frontier inequality in $\mathcal{K}^{\prime}$ we have that
@@
First, we have to define a notion of smoothness meeting Definition 3.27. We declare the smooth locus of a definable set $X$ to be the set of points $x\in X$ where $X$ is, locally near $x$, a $\dim(X)$-dimensional weak strictly differentiable manifold (see [ALH25, Definition 5.3]). By visceral cell decomposition, every definable set is generically locally a topological manifold, and by Proposition 3.12 loc. cit. it is, in fact, generically, a strictly differentiable manifold. This gives all the properties of Definition 3.27 except maybe (5), the preservation of smoothness under preimages of “nice enough” projections. But recall (see the remark following Definition 3.27) that this clause is a generalization of, and would follow from, the ‘submersion theorem’ of differential geometry. Moreover, there is a submersion theorem for 1-h-minimal field ([ALH25, Proposition 4.10]). So we simply apply that proposition exactly as in the earlier remark.
@@
We use the usual notion of a tangent space associated with the notion of (weak) strictly differentiable manifolds ([ALH25, Definition 5.4]). All properties (1)-(7) of the definition are automatic, and we don’t dwell on them. Clause (8) of the definition, asserting that a definable morphism inducing an isomorphism on the level of tangent spaces is open (locally, near the point in question) is an immediate consequence of the Inverse Mapping Theorem ([ALH25, Proposition 4.4]). So it only remains to check that if $f:(X,x)\to(Y,y)$ is such that $x\in X$ and $y\in Y$ are generic, then the induced map between the tangent spaces is surjective. This, in turn, is an immediate result of the corresponding version of Sard’s Theorem [ALH25, Proposition 5.21], asserting that the set of singular values of $f$ is non-generic.